# Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era

### v1

@@
Abstract
Explainable AI (XAI) refers to techniques that provide human-understandable insights into the workings of AI models. Recently, the focus of XAI is being extended towards Large Language Models (LLMs) which are often criticized for their lack of transparency. This extension calls for a significant transformation in XAI methodologies because of two reasons. First, many existing XAI methods cannot be directly applied to LLMs due to their complexity and advanced capabilities. Second, as LLMs are increasingly deployed across diverse industry applications, the role of XAI shifts from merely opening the “black box” to actively enhancing the productivity and applicability of LLMs in real-world settings. Meanwhile, unlike traditional machine learning models that are passive recipients of XAI insights, the distinct abilities of LLMs can reciprocally enhance XAI. Therefore, in this paper, we introduce Usable XAI in the context of LLMs by analyzing (1) how XAI can benefit LLMs and AI systems, and (2) how LLMs can contribute to the advancement of XAI. We introduce 10 strategies, introducing the key techniques for each and discussing their associated challenges. We also provide case studies to demonstrate how to obtain and leverage explanations. The code used in this paper can be found at: https://github.com/JacksonWuxs/UsableXAI_LLM.
**footnotetext: Equal contribution
1 Introduction
Explainability holds great promise in understanding machine learning models and providing directions for improvement. In practice, users have high expectations for model explainability:
@@ 1. Through explanation, can we know if a model works properly?
2. Does explainability help developing better models?
@@
First, explanations are expected to illuminate whether a model operates in accordance with human expectations. For example, does the model leverage reliable evidence and domain knowledge in its decision-making? Does the model contain bias and discrimination? Does the model show any vulnerabilities to potential attacks? Will the model output harmful information? Second, in recognition of model imperfections, we aspire for explainability to inform the development of better models. For example, how to adjust the behaviors of a model if we find it is using unreliable or unreasonable features in making predictions? Can we improve the performance of a model by aligning its behavior with human preferences?
@@
Therefore, the question arises: Have these expectations been met? In recent years, the body of literature on Explainable AI (XAI) has expanded rapidly to improve model transparency (Du et al., 2019a; Murdoch et al., 2019; Tjoa & Guan, 2020; Došilović et al., 2018; Rudin et al., 2022), encompassing a wide array of methods customized for different data modalities, including visual (Zhang & Zhu, 2018), textual (Danilevsky et al., 2020), graph (Yuan et al., 2022), and time-series data (Zhao et al., 2023c). Some literature delves into specific techniques, such as attention methods, generalized additive models, and causal models. Additionally, some offer reviews on general principles and categorizations or initiate discussions on evaluating the faithfulness of explanations (Yang et al., 2019). Despite the progress, the last mile of XAI – making use of explanations – has not received enough attention. In many cases, we seem to be satisfied with just acquiring explanations and their associated visualizations, sometimes followed by qualitative analysis of the model’s strengths and weaknesses. While these explanations can reveal a model’s imperfections, quantifying model properties (e.g., fairness, security, rationality) or taking the next concrete steps toward model improvement remains a difficult task.
The challenges in achieving usable explainability are twofold. First, there is an inherent conflict between AI automation and human engagement in XAI. On one hand, humans need to define explainability that the model should follow, or scrutinize explanation to identify if any vulnerabilities exist in the model. On the other hand, the requirement for human oversight introduces substantial costs, posing challenges to the scalability and practical implementation of model debugging and improvement in AI workflows. Second, many of the current approaches view explainability as a purely technical matter, ignoring the needs of practitioners and non-technical stakeholders. Existing XAI methods are mainly developed as statistical and mathematical tools. However, there exists a noticeable disparity between the objectives of these tools and the expectations of practitioners across various application domains (Malizia & Paternò, 2023). An explanation that satisfies a technical audience might offer little value to a non-technical audience.
@@
While the opacity issues have not yet been fully resolved for traditional deep models (e.g., multi-layer perceptrons, convolutional and recurrent neural network), the recent advancements of Large Language Models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023b; Chiang et al., 2023) appear to have exacerbated the challenge we are facing. Firstly, LLMs typically possess a significantly larger model size and a greater number of parameters. This increased model complexity intensifies the difficulty of explaining their inner workings. Second, different from traditional ML models that primarily focus on low-level pattern recognition tasks such as classification and parsing, LLMs can handle more complex tasks such as generation, reasoning and question answering. Understanding the exclusive abilities of LLMs presents novel challenges for XAI techniques. Considering the transformative impact of LLMs across various applications, ensuring the explainability and ethical use of LLMs has become an imminent and pressing need. Meanwhile, the emergent capabilities of LLMs also present new opportunities for XAI research. Their human-like communication and commonsense reasoning skills offer prospects for achieving explainability in ways that could potentially augment or replace human involvement.
@@
Figure 1: The contributions and outline of this paper. We define Usable XAI in the context of LLMs with seven strategies of enhancing LLMs with XAI, and three strategies of enhancing XAI with LLMs.
@@
Defining “Usable XAI”. In light of the above considerations, in the context of LLMs, we define Usable XAI which includes two aspects as follows. (1) Utilizing Explainability to Enhance LLM and AI Systems. Beyond just producing explanations or enhancing the transparency of LLMs, we explore whether these explanations can pinpoint issues for model debugging or improve the overall performance of LLMs or AI models at large, such as accuracy, controllability, fairness, and truthfulness. (2) Utilizing LLMs to Enhance XAI Frameworks. The human-like communication ability of LLMs can enhance model explanations in terms of user-friendliness, by converting the numerical values into understandable language. Also, the commonsense knowledge stored in LLMs can significantly boost the practicality of existing XAI frameworks, by playing the role of humans and alleviating the need for real human involvement in AI workflows.
@@
Contribution of this paper. In this paper, we investigate 10 strategies towards usable XAI techniques in the context of LLMs. These strategies are organized into two major categories: (1) Usable XAI for LLMs; (2) LLM for Usable XAI, as shown in Figure 1. Additionally, we conduct case studies to substantiate the discussion on selected techniques. For each strategy, we also explore the open challenges and areas that require further investigation in future work.
@@
  • Usable XAI for LLMs. We introduce how interpretation can be utilized to enhance AI pipelines, including LLMs and small models. First, we investigate how explanations could be utilized to diagnose and enhance LLMs in terms of utility. We study three types of post-hoc explanation methods, targeting LLM predictions (Section 2), LLM components (Section 3), and training samples (Section 4), respectively. Second, we focus on how explanations could be leveraged to scrutinize and boost model trustworthiness (Section 5), including security, fairness, toxicity, and truthfulness, which is crucial to achieving human alignment. Third, we discuss how explainability could guide the augmentation of data, including both inference data (i.e., prompts) and training data. Specifically, we discuss two strategies of crafting explainable prompts for LLMs: Chain-of-Thought prompts (Section 6) and knowledge-enhanced prompts (Section 7). Furthermore, we introduce leveraging LLM explanations to augment training data for improving small models (Section 8).
@@
  • LLM for Usable XAI. In this part, we investigate strategies for leveraging the advanced capabilities of LLMs to address the challenges in traditional XAI domains, thus enhancing the usability of XAI in practice. First, we examine ways to enhance the user-friendliness of explanations through the generative capabilities of LLMs (Section 9). Second, we introduce how to automate the design of interpretable AI workflows by leveraging the planning abilities of LLMs (Section 10). Third, we introduce how to facilitate the evaluation of XAI methods by utilizing the unique property of LLMs in emulating human cognition processes (Section 11).
@@
Differences between this paper and existing surveys. Many surveys have been conducted to examine Explainable AI (Du et al., 2019a; Tjoa & Guan, 2020; Došilović et al., 2018) or Interpretable Machine Learning (Murdoch et al., 2019). This paper differs from existing work as we focus on explanation methods for large language models. Meanwhile, different from the existing survey (Zhao et al., 2023b) that mainly reviews explanation methods for LLMs, our paper puts an emphasis on the XAI usability in LLM studies. To the best of our knowledge, the most related paper to our survey is (Luo & Specia, 2024), which also discusses several aspects where explanations can improve LLM performance. Nevertheless, this light-weight investigation lacks a thorough examination of XAI methods (e.g., sample-based explanation, interpretable workflows, explainable prompts) and how LLMs can benefit existing XAI frameworks (e.g., data augmentation, improving user-friendliness, XAI evaluation). Finally, our paper contributes further by providing detailed case studies and open-sourced codes, fostering future research in applying explanations effectively within the LLM context.
2 LLM Diagnosis via Attribution Methods
@@
This section introduces attribution methods as post-hoc explanations for LLMs, and how we can discover model defects with attribution scores. We start with revisiting existing attribution methods, and then discuss which methods are still suitable for explaining LLMs. Since LLMs widely serve both classification and generation tasks, our discussion categorizes the attribution methods accordingly. After that, we explore case studies of applying attribution methods to assess LLM-generated output quality. Finally, we discuss future work of designing novel post-hoc explanation methods for LLMs.
2.1 Literature Review of Attribution Methods
@@
The attribution-based explanation quantifies the importance of each input feature that contributes to making predictions. Given a language model $f$ with a prediction $\hat{y}=f(x)$ according to the $N$-words input prompt $x$, the explainer $g$ assesses the influence of input words in $x$ as $\mathbf{a}=g(x,\hat{y},f)\in\mathbb{R}^{N}$. Typically, the sign of $a_{n}\in\mathbf{a}$ indicates word $x_{n}$ positively or negatively influences $\hat{y}$, and a greater value of $|a_{n}|$ indicates a stronger impact. In text classification, $\hat{y}$ denotes a specific class label. In text generation, $\hat{y}$ represents a varying length of generated text.
Many existing attribution-based explanation methods focus on classification tasks and cannot be directly applied to the generation task. The primary distinction between them is that: classification is limited to a specific set of predictions, while generation encompasses an endless array of possibilities. For instance, in sentiment analysis, a language model can be instructed to output a number between 0 and 1 that indicates the positivity of input text by adding a linear layer and a sigmoid function on top of the language model. However, in the generative setting, the model can express this positivity in numerous expressions, such as “the reviewer definitely loves this movie” and “it is a strong positive movie review”. This distinction poses a unique challenge in adapting explanation methods from classification to generation tasks. In the following, we review related works based on the scenarios they are applicable to.
2.1.1 Attributing Inputs for Label Classification
@@
Common attribution methods (Du et al., 2019a; Murdoch et al., 2019) developed for traditional deep models include gradient-based methods, perturbation-based methods, surrogate methods, and decomposition methods. We introduce the general idea and representative examples for each category, followed by the analysis of their suitability for explaining large language models.
@@
Perturbation-based Explanation. Perturbation-based methods assess the importance of input features by perturbing them and monitoring changes in prediction confidence, i.e., $a_{n}=p(\hat{y}|x)-p(\hat{y}|\widetilde{x}_{n})$, where $\widetilde{x}_{n}$ refers to the input sequence with the $n$-th feature being perturbed. Each feature could refer to a word (Li et al., 2016a), a phrase (Wu et al., 2020b), or a word embedding (Li et al., 2016b). The underlying principle is that perturbing a more important feature should result in a more pronounced alteration in the model’s prediction confidence. However, this method has limitations, particularly in its assumption that features are independent, which is not always the case with textual data due to word inter-dependencies. Additionally, it is computationally intensive for explaining LLMs, requiring $N$ inferences for an input of $N$ words.
Gradient-based Explanation. Gradient-based methods offer a computationally efficient approach for estimating model sensitivity to input features based on gradients $\frac{\partial p(\hat{y}|x)}{\partial\mathbf{x}_{n}}$, where $\mathbf{x}_{n}$ refers to the embedding of word $x_{n}$. Some methods employ the $L_{2}$-norm of gradients to assess word importance (Li et al., 2016a), i.e., $a_{n}=\|\frac{\partial p(\hat{y}|x)}{\partial\mathbf{x}_{n}}\|_{2}$. This approach only requires a single inference and one backpropagation pass. Some extended methods multiply the gradient with the word embedding (Kindermans et al., ; Ebrahimi et al., 2018; Mohebbi et al., 2021), i.e., $a_{n}=\frac{\partial p(\hat{y}|x)}{\partial\mathbf{x}_{n}}\cdot\mathbf{x}_{n}$. These methods may yield explanations with limited faithfulness for deep models (Shrikumar et al., 2017), as gradients only reflect the local relationship between input variation and output variation. To address this, Integrated Gradients (IG) has been proposed (Sundararajan et al., 2017; Sikdar et al., 2021; Sanyal & Ren, 2021; Enguehard, 2023), which accumulates gradients as input transitions from a reference point to the actual data point. Nevertheless, IG entails multiple rounds of inference and backpropagation, thus significantly increasing computational demands.
@@
Surrogate-based Explanation. Surrogate-based explanation methods understand complex models by constructing a simpler model $g$ trained on $\mathcal{D}(x,\hat{y})=\{(\widetilde{x}_{k},\widetilde{y}_{k})\}_{k=1}^{K}$, where $\mathcal{D}(x,\hat{y})$ denotes a dataset constructed for the target instance $(x,\hat{y})$; $\widetilde{x}_{k}$ is usually obtained by perturbing $x$, and $\widetilde{y}_{k}=f(\widetilde{x}_{k})$. The surrogate model $g$, ranging from basic linear models to sophisticated decision trees, serves as a proxy to approximate the decision boundary of the target model $f$ for a specific instance $(x,\hat{y})$. Notable examples include LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017), and TransSHAP (Kokalj et al., 2021), where the first two are designed for general deep neural networks and the last one is tailored for Transformer-based language models. Nevertheless, a significant limitation of them is their intensive reliance on repeated interactions with the target model, a process that is impractical for LLMs.
@@
Decomposition-based Explanation. Decomposition-based methods assign linearly additive relevance scores to inputs, effectively breaking down the model’s prediction. Layer-wise Relevance Propagation (Montavon et al., 2019) and Taylor-type Decomposition (Montavon et al., 2017) are well-known techniques for computing these relevance scores. These methods have been adapted for Transformer-based language models in various research (Voita et al., 2019; 2020; Wu & Ong, 2021). However, a primary challenge in implementing decomposition-based explanations is the need for tailored decomposition strategies to accommodate different model architectures. Although many large language models are based on the Transformer framework, there are key variations between them, such as LLaMA (Touvron et al., 2023a) and GPT (OpenAI, 2023), particularly in aspects like positional encoding strategy and feed-forward network design. This challenge poses a limitation on the universal applicability of decomposition methods for general-purpose interpretation.

### v2

@@
Abstract.
Explainable AI (XAI) refers to techniques that provide human-understandable insights into the workings of AI models. Recently, the focus of XAI is being extended toward explaining Large Language Models (LLMs). This extension calls for a significant transformation in the XAI methodologies for two reasons. First, many existing XAI methods cannot be directly applied to LLMs due to their complexity and advanced capabilities. Second, as LLMs are increasingly deployed in diverse applications, the role of XAI shifts from merely opening the “black box” to actively enhancing the productivity and applicability of LLMs in real-world settings. Meanwhile, the conversation and generation abilities of LLMs can reciprocally enhance XAI. Therefore, in this paper, we introduce Usable XAI in the context of LLMs by analyzing (1) how XAI can explain and improve LLM-based AI systems and (2) how XAI techniques can be improved by using LLMs. We introduce 10 strategies, introducing the key techniques for each and discussing their associated challenges. We also provide case studies to demonstrate how to obtain and leverage explanations. The code used in this paper can be found at: https://github.com/JacksonWuxs/UsableXAI_LLM.
Large Language Models, Explainable AI (XAI), Usability
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Machine learning
1. Introduction
Explainability holds great promise for understanding machine learning models and providing directions for improvement. In practice, users have high expectations for model explainability:
@@ 1. Through explanation, can we know if a model works properly?
2. Does explainability tell us how to develop better models?
@@
First, explanations are expected to illuminate whether a model operates as humans expect. For example, does the model leverage reliable evidence and domain knowledge in its decision making? Does the model contain bias and discrimination? Does the model show any vulnerabilities to potential attacks? Will the model output harmful information? Second, in recognition of model imperfections, we aspire for explainability to inform the development of better models. For example, how to adjust the behaviors of a model if we find that it uses unreliable or unreasonable features in making predictions? Can we improve the performance of a model by aligning its behavior with human preferences?
@@
Therefore, the question arises: Have these expectations been met? Since the rise of deep learning, the body of literature on Explainable AI (XAI) has expanded rapidly to improve model transparency (Du et al., 2019a; Murdoch et al., 2019; Tjoa and Guan, 2020; Došilović et al., 2018; Rudin et al., 2022), encompassing a wide array of methods customized for different data modalities, including visual (Zhang and Zhu, 2018), textual (Danilevsky et al., 2020), graph (Yuan et al., 2022), and time-series data (Zhao et al., 2023c). In addition, some offer reviews of general principles and initiate discussions on evaluating the faithfulness of explanations (Yang et al., 2019; Doshi-Velez and Kim, 2017). Despite the progress, the last mile of XAI – making use of explanations – has not received the attention it merits. In many cases, we seem to be satisfied with just acquiring explanations and their associated visualizations, sometimes followed by qualitative discussions of the model’s strengths and failure cases. However, how to quantify model properties (e.g., fairness, security, rationality) and improve models by utilizing explanations remains a difficult task.
@@
While the opacity issues have not yet been fully resolved for traditional deep models (e.g., multi-layer perceptrons, convolutional and recurrent neural networks), the recent advancements of Large Language Models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023b; Chiang et al., 2023) appear to have exacerbated the challenge we are facing. Firstly, LLMs typically possess a significantly larger model size and a greater number of parameters. This increased model complexity intensifies the difficulty of explaining their inner workings. Second, unlike traditional ML models that primarily focus on low-level pattern recognition tasks such as classification and parsing, LLMs can handle more complex tasks (Zhao et al., 2023d) such as generation, reasoning, and question answering. Understanding the exclusive abilities of LLMs presents novel challenges for XAI techniques. Considering the transformative impact of LLMs across various applications, ensuring the explainability and ethical use of LLMs has become an imminent and pressing need. Meanwhile, the emerging capabilities of LLMs also present new opportunities for XAI research. Their human-like communication and commonsense reasoning skills offer prospects for achieving explainability in ways that could potentially augment or replace human involvement.
@@
Figure 1. The contributions and outline of this paper. We define Usable XAI in the context of LLMs with seven strategies of enhancing LLMs with XAI, and three strategies of enhancing XAI with LLMs.
@@
Defining “Usable XAI”. In light of the above considerations, in the context of LLMs, we define Usable XAI which includes two aspects. (1) Leveraging XAI to improve LLMs and AI Systems. Beyond just producing explanations of LLMs, we explore whether these explanations can pinpoint issues for model debugging or improve the overall performance of LLMs or AI models, such as accuracy, controllability, and trustworthiness. (2) Enhancing the usability of XAI through LLMs. The human-like communication ability of LLMs can enhance model explanations in terms of user-friendliness, by converting the numerical values into understandable language. Also, the commonsense knowledge stored in LLMs can significantly boost the practicality of existing XAI frameworks, by playing the role of humans and alleviating the need for real human involvement in AI workflows.
@@
Contribution of this paper. In this paper, we investigate 10 strategies towards usable XAI techniques in the context of LLMs, as shown in Figure 1. For each strategy, we also explore the open challenges that require further investigation in future work.
@@
  • Leveraging XAI to improve LLMs and AI Systems We introduce how interpretation can be utilized to enhance LLMs and AI models. We follow the LLM development pipeline in reverse order—spanning post-hoc analysis, prompt improvement, and data enhancement. First, we investigate how post-hoc explanations could be utilized to diagnose and enhance LLMs. We study three types of post-hoc explanation methods, targeting LLM predictions (Section 2), LLM components (Section 3), and training samples (Section 4), respectively, followed by how explanations could be leveraged to scrutinize and boost LLM trustworthiness (Section 5), including security, fairness, toxicity, and truthfulness. Second, we discuss two strategies for designing explainable prompts to guide LLM reasoning: Chain-of-Thought prompts (Section 6) and knowledge-enhanced prompts (Section 7). Third, we introduce the use of LLM explanations to augment training data toward improving AI models (Section 8).
@@
  • Enhancing XAI Usability through LLMs In this part, we investigate strategies for leveraging the advanced capabilities of LLMs to address the challenges in traditional XAI domains, thus enhancing the usability of XAI in practice. First, we examine ways to enhance the user-friendliness of explanations through the generative capabilities of LLMs (Section 9). Second, we introduce how to automate the design of interpretable AI workflows by leveraging the planning abilities of LLMs (Section 10). Third, we introduce how to facilitate the evaluation of XAI methods by utilizing the unique property of LLMs in emulating human cognition processes (Section 11).
@@
Differences between this paper and existing surveys. Many surveys have been conducted to examine Explainable AI (Du et al., 2019a; Tjoa and Guan, 2020; Došilović et al., 2018; Murdoch et al., 2019). This paper differs from existing work as we focus on explanation methods for large language models. Meanwhile, different from the existing survey (Zhao et al., 2023a; Ferrando et al., 2024) that reviews explanation methods for LLMs, our paper places an emphasis on the XAI usability in LLM studies. To the best of our knowledge, the most related paper to our survey is (Luo and Specia, 2024), which also discusses several aspects where explanations can improve LLM performance. Nevertheless, this light-weight investigation lacks a thorough examination of XAI methods (e.g., sample-based explanation, interpretable workflows, explainable prompts) and how LLMs can benefit existing XAI frameworks (e.g., data augmentation, improving user-friendliness, XAI evaluation). Finally, our paper contributes further by providing detailed case studies and open-sourced codes, fostering future research in applying explanations effectively within the LLM context.
2. LLM Diagnosis and Enhancement via Attribution Methods
@@
This section introduces attribution methods as post-hoc explanations for LLMs, and how we can discover model defects with attribution scores. We start by reviewing existing attribution techniques and discussing which methods are still suitable for explaining LLMs. After that, we explore cases that apply attribution methods to assess LLM-generated output quality. Finally, we discuss future work of designing novel post-hoc explanations for LLMs.
2.1. Attribution Methods for LLMs
@@
The attribution-based explanation quantifies the importance of each input feature that contributes to the prediction. Given a language model $f$ with a prediction $\hat{y}=f(x)$ according to the $N$-words input prompt $x$, the explainer $g$ assesses the influence of input words in $x$ as: $a=g(x,\hat{y},f)$, where $a$ stores the importance scores of words. In text classification, $\hat{y}$ denotes a specific class label. In text generation, $\hat{y}$ represents a varying length of generated text. The primary distinction between them is that: classification is limited to a specific set of predictions, while generation encompasses an endless array of possibilities. In the following, we review related works based on the tasks to which they are applicable.
2.1.1. Attribution Methods for Classification
@@
Common attribution methods (Du et al., 2019a; Murdoch et al., 2019) for deep models include gradient-based methods, perturbation-based methods, surrogate methods, and decomposition methods.
@@
  • Perturbation-based explanation assesses the importance of input features by perturbing them and monitoring changes in prediction confidence, i.e., $a_{n}=p(\hat{y}|x)-p(\hat{y}|x_{/n})$, where $x_{/n}$ refers to the input sequence with the $n$-th feature being perturbed or deleted (Li et al., 2016a, b; Wu et al., 2020b). This approach has limitations, particularly in its assumption that features are independent, which is not always the case with textual data. Additionally, it is computationally intensive for explaining LLMs, requiring $N$ inferences for an input of $N$ words.
  • Gradient-based explanation offers a computationally efficient approach for estimating model sensitivity to input features based on gradients $\frac{\partial p(\hat{y}|x)}{\partial\mathbf{x}_{n}}$, where $\mathbf{x}_{n}$ refers to the embedding of word $x_{n}$ (Li et al., 2016a; Kindermans et al., [n. d.]; Shrikumar et al., 2017; Ebrahimi et al., 2018; Mohebbi et al., 2021). This approach only requires a single inference and one backpropagation pass, achieving an efficient way to estimate the input word sensitivities. However, it is still expensive to compute a backward pass on such giant LLMs.
@@
  • Surrogate-based explanation understands the target model $f$ by constructing a simpler model $g$ trained on $\mathcal{D}(x,\hat{y})=\{(\widetilde{x}_{k},\widetilde{y}_{k})\}_{k=1}^{K}$, where $\widetilde{x}_{k}$ is usually obtained by perturbing $x$, and $\widetilde{y}_{k}=f(\widetilde{x}_{k})$ (Ribeiro et al., 2016; Lundberg and Lee, 2017; Kokalj et al., 2021). The surrogate model $g$, ranging from basic linear models to sophisticated decision trees, serves as a proxy to approximate the decision boundary of $f$ for the target instance $(x,\hat{y})$. Nevertheless, a significant limitation of this approach is its repeated interactions with the target model, which leads to huge computing requirements. In addition, it is impractical to find a simple and explainable model $g$ that can estimate the advanced LLMs well.
@@
  • Decomposition-based explanation assigns linearly additive relevance scores to inputs, effectively breaking down the model’s prediction (Montavon et al., 2019, 2017; Voita et al., 2019, 2020; Wu and Ong, 2021). However, a primary challenge in implementing such an approach is the need for tailored decomposition strategies to accommodate different model architectures. This challenge poses a limitation on the universal applicability of decomposition methods for general-purpose interpretation.
2.1.2. Attribution Methods for Generation
With advancements in generative AI, interpreting why an output is generated has become increasingly important. The explanation of LLM generation can be defined as attributing the overall confidence $p(\hat{y}|x)$ to the input $x$, where $\hat{y}$ denotes the generated response $\hat{y}=[\hat{y}_{1},...,\hat{y}_{M}]$ with $M$ words. An approach is to treat the text generation process as a sequence of word-level classification tasks. This perspective allows for the application of existing classification-based explanation techniques to assess the influence of each input word $x_{n}$ in relation to each output word $\hat{y}_{m}$. In particular, Wu et al. (2023b) proposes a gradient-based method for input-output attribution. Specifically, the importance score $a_{n,m}$ defined between input token $x_{n}$ to output token $\hat{y}_{m}$ can be estimated as:
(1) $a_{n,m}=p(\hat{y}_{m}|x,\hat{y}_{1:(m-1)})-p(\hat{y}_{m}|x_{/n},\hat{y}_{1:(m-1)})\approx\frac{\partial f(\hat{y}_{m}|x,\hat{y}_{1:(m-1)})}{\partial\mathbf{E}[x_{n}]}\cdot\mathbf{E}[x_{n}],$
where $\hat{y}_{1:(m-1)}$ denotes the first $m-1$ tokens of response, $x_{/n}$ means removing the $n$-th input token from $x$, and $\mathbf{E}[x_{n}]$ is the embedding of token $x_{n}$. The total contribution of each input word $x_{n}$ can be estimated by averaging the attributions $a_{n,m}$ for $m=1,...,M$ (Selvaraju et al., 2016). Figure 2 (left) plots an example, where each index in the Y-axis refers to an input token, while that in the X-axis is an output token. A higher attribution score is brighter. Given that not all output tokens warrant interpretation, Qi et al. (2024) outlines a two-step procedure where informative tokens are first identified, and then attribution is made to these tokens using gradient-based or other attribution techniques (Yin and Neubig, 2022). It also develops a contrastive explanation method, which aims to explain why the model predicted one target token $\hat{y}_{m}$ over an alternative token $\hat{y}^{\prime}_{m}$. Furthermore, a Python library has been developed to incorporate existing attribution methods for sequential generation of language models (Sarti et al., 2023).
Time complexity is a critical consideration when applying attribution methods to language generation models like LLMs. In particular, the perturbation-based and gradient-based approaches may be limited by their huge demand for computing resources, while the surrogate-based and decomposition-based strategies are restricted by their limited generalizability across different model architectures. To emphasize the computing costs of applying attribution-based approaches for generative LLMs, we present the time complexity of several representative methods in Table 1, by assuming the model takes $N$ words as input and predicts $M$ words as output. It shows that existing attribution methods require a large number of forward or backward operations, emphasizing the exploration of efficient methods.